OpenStudy (anonymous):

Find the distance for the following two points A (-10,-15) (-5, -3). 1) 13 2) -13 3) 12 4) -12

OpenStudy (anonymous):

-5 - (-10) = 5 -3 - (-15) = 12 sqroot ( 5^2 + 12^2 ) = 13

OpenStudy (lgbasallote):

\[distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[distance = \sqrt{[-5 - (-10)]^2 + [(-3 - (-15)]^2}\] \[distance = \sqrt{(-5+10)^2 + (-3 + 15)^2}\] \[distance = \sqrt{(-5)^2 + (12)^2}\] \[distance = \sqrt{25 + 144} \] \[distance = \sqrt{169}\] \[distance = 13\]

OpenStudy (anonymous):

In the fourth step is positive 5 even do it doesnt affect the final result